OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y = one over twox - 3 using the slope-intercept method? Plot the y-intercept (0, -3). Draw a line through the two points. Count up 1 and right 2 from the y-intercept. Plot the x-intercept (6, 0).

OpenStudy (rose21):

x - 2y = 4 ..it becomes.. -2y = -x + 4 ..which becomes.. y = 1/2x -2 Your slope is 1/2 (0.5) and your y-intercept is -2

OpenStudy (anonymous):

thank you is the answer A than

OpenStudy (anonymous):

@rose21

OpenStudy (rose21):

I think so